# Decoupled-Value Attention for Prior-Data Fitted Networks: GP Inference for Physical Equations

### v1

@@ 2.3 Limitations of Existing PFN Architectures with Joint Attentions
PFNs offer a promising framework for amortized Bayesian inference, though their application to high-dimensional regression has so far been relatively limited Hollmann et al. (2025); Wang et al. (2025). The common recipe of using a Transformer backbone that performs self-attention over joint $(\mathbf{x},\mathbf{y})$ embeddings has a scaling issue. The design choice of representing each training example $(x,y)$ in PFNs as a joint embedding $\mathrm{enc}(x)+\mathrm{enc}(y)$ can be traced back to the way attention-based models have historically treated their basic units of computation. In natural language processing, the Transformer architecture Vaswani et al. (2017) encodes each token as a self-contained representation, i.e. “token as full carrier of information”. In this lineage, PFNs adopt the same strategy by concatenating or summing input and output encodings to form a single token while removing positional encoding Müller et al. (2022). Below, we examine the this PFN recipe’s structural limitations.
@@
Further, we can also analyze this dimensionality limitation from a bias perspective as the joint embedding breaks localization. The Transformer computes similarity (via dot-product attention) between queries and these mixed embeddings in which the label $\mathbf{y}_{i}$ contributes equally. This conflicts with theoretical results Nagler (2023), which show that only local samples should influence posterior estimates. Consequently, incorporating joint input–output attention introduces additional bias, which becomes more pronounced in higher dimensions due to the concentration of pairwise distances. In view of these limitations, we propose a simple decoupled value attention (DVA) which keeps the localization intact.
@@ 3 Decoupled-Value Attention
We propose Decoupled-Value Attention (DVA), an input-localized attention mechanism for training PFNs. The proposed DVA is structurally aligned with GP inference by treating input $\mathbf{x}$ and output $\mathbf{y}$ separately at the attention stage. We enforce a strict separation of roles: attention affinities (queries and keys) are computed solely from the inputs, while the aggregated information (values) comes from the corresponding outputs– during both PFN training and prediction. Below, we explain DVA mathematically along with comparative assessment against Vanilla Attention (VA) Müller et al. (2022) and a kernel-based attention Wang & Others (2025).
@@
Training: During training we simulate inference by masking one or more labels from the dataset Müller et al. (2022). The unmasked pairs $\mathcal{D}_{\mathrm{cx}}=\{(x_{i},y_{i})\}_{i=1}^{N_{\mathrm{context}}}$ form the context set, while the masked inputs $X_{\mathrm{te}}=\{x_{j}\}_{j=1}^{M}$ form the queries. From the context we build
$K_{\mathrm{tr}}=W_{k}\varphi_{x}(X_{\mathrm{cx}}),\quad V_{\mathrm{tr}}=W_{v}\varphi_{y}(\mathbf{y}_{\mathrm{cx}}),$ (3)
where, $X_{\mathrm{cx}}$ and $\mathbf{y}_{\mathrm{cx}}$ are matrix and vector forms of $\mathcal{D}_{\mathrm{cx}}$ respectively. Further, from the test (masked) inputs query $Q_{\mathrm{te}}$ in matrix from and labels are predicted by attending $H_{\mathrm{te}}$ to the context as :
@@
A head $g(\cdot)$ maps $H_{\mathrm{te}}$ to a predictive distribution, and training minimizes the NLL (equation 1) of the true labels to learn parameters of the network as explained in Müller et al. (2022).
@@
Vanilla attention is taken from PFN literature Müller et al. (2022); Hollmann et al. (2022)
@@ 3.1 Localization Effect of DVA and Alignment with GP inference
From equation 6, it is clear that that affinities are determined entirely via relationship between the test input $\mathbf{x}_{\star}$ and context inputs $X_{i}$. Unlike joint embeddings $\phi(x,y)$ in VA, the labels $y_{i}$ do not enter into the similarity measure and only appear downstream through the values as in equation 2. This separation implies that the weight placed on a output $y_{i}$ depends solely on how well $X_{i}$ aligns with $\mathbf{x}_{\star}$ in the projected input space. Thus, the softmax distribution $\alpha_{i}(\mathbf{x}_{\star})$ concentrates mass on a neighborhood of $\mathbf{x}_{\star}$ because the projection matrices and encoders are trained to align nearby inputs with high value of inner product and push apart distant inputs. Consequently, altering labels (outputs) attached to distant inputs cannot affect the prediction asymptotically, which is exactly the localization property required in Theorem 5.4 of Nagler (2023). Thus, proposed DVA recovers input-space localization by construction, while still using the standard softmax form of attention.
We now discuss how DVA aligns with GP inference. As discussed in Section 2.1, GPs model all possible function realizations as zero-mean Gaussians with covariance defined by a kernel, i.e., $f\sim\mathcal{N}(\mathbf{0},K(X,X^{\prime}))$, where $X$ is the input and $K(\cdot,\cdot)$ is the kernel matrix over input pairs. Note that parameterization of the possible function family only depends on the input. Further, for a given kernel hyperparameters, the mean prediction $\mu(\cdot)$ of GP is given as a weighted sum of training dataset outputs with weights solely depending upon inputs Williams & Rasmussen (2006):
@@
Following equation 5, 6 and 7, the attention weights in DVA can be interpreted as normalized kernel weights that depend only on the inputs. As in kernel smoothing Tsai et al. (2019), the exponential inner product in $\alpha(\cdot)$ of equation 6 acts as a positive kernel on the input space, with effective bandwidth governed by the scale of the projections and the $1/\sqrt{d_{k}}$ factor. Thus, similar to GP mean prediction, DVA predictions are obtained as weighted sums of training outputs where the weights are determined entirely by input similarity. Readers can refer to Tsai et al. (2019) for more discussion on the relationship between kernel and attention mechanism.
Here, we want to highlight that DVA’s softmax produces non-negative, normalized weights ($\sum\alpha_{i}=1$), whereas the GP coefficients $\beta_{i}(\cdot)$ have no positivity constraint. This limitation is mitigated by subsequent PFN layers (e.g., the final head $g(\cdot)$) and by encoding outputs in the value $V$, which together help adjust the DVA output toward the true GP posterior mean. This construction shows that DVA’s architecture implements a predictor of the form “input-only similarities produce weights, which combine label-dependent values,” precisely matching the dependency structure of a GP.
Another attention choice for PFNs can be kernel-inspired attentions, which relate GP mean weights $\beta(\cdot)$ and PFN attention weights $\alpha(\cdot)$ more closely– while maintaining input localization by decoupling input and output as in DVA. However, if the input affinities are forced through a fixed kernel function, the PFN will become kernel dependent. As discussed before, identifying the best performing kernel is non-trivial and often requires tailoring kernels to specific function classes Liu & Srikantha (2022). Therefore, it is not advisable to hard-wire a particular kernel in PFN design.
To test the effect of kernel dependence on PFN performance, we design a simple Gaussian kernel (radial basis function, RBF) similarity for attention Williams & Rasmussen (2006). We emphasize that this formulation is not equivalent to exact GP kernel regression but rather introduces RBF-style distance-based affinities in place of dot-product similarities Choromanski et al. (2022); Shen et al. (2021). The kernel-based attention assigned to a query input $\mathbf{x}_{\star}$ is then given by
@@
This distance-based attention in equation 8 is more aligned to the RBF kernel; however, it loses flexibility to learn input-localization via training. Thus, the model inherits kernel and $\gamma$ dependence, which may not be suitable for a broader class of functions. To validate this limitation of the kernel-based attention, we test PFN performance with both DVA and this attention. We attempt to learn two classes of functions with different levels of smoothness as shown in Figure 1. The results demonstrate that while kernel-based attention can match DVA in effectively learning smooth functions aligned with the RBF kernel, it significantly underperforms on non-smooth functions generated using the linear-periodic kernel. More details on this experiment can be found in Appendix C.
@@ 4 Numerical Results and Discussion
In this section, we present numerical experiments demonstrating the behavior of PFNs equipped with the proposed DVA and with CNN backbone. The results show that with DVA, PFNs a) train with lower validation loss or residual bias, b) both CNN and Transformer perform comparably as architecture, underscoring that attention governs training behavior more than backbone architecture and c) remain scalable for learning in complex physical systems. These findings provide empirical support for the theoretical arguments in Section 3.1. Complete experimental details, including architecture choices, hyperparameter selection, and data generation procedures, are provided in the Appendix B, while additional results are provided in Appendix D.
4.1 Bias Reduction and CNN PFNs
@@
Figure 2: Bias Reduction in PFN Training: Validation loss (NLL) behavior with number of training points for various PFNs (Number of training points = epochs $\times$ steps per epoch $\times$ batch-size $\times$ dataset size. Dataset size is 100 for 1D/2D, 400 for 5D and 500 for 10D PFN). Validation loss was calculated on 64 out-of-sample datasets and Transformer + VA is taken from Müller et al. (2022).
Comparative Discussion on CNN- and Transformer-Based PFNs: To analyze the effect of backbone architecture on PFN performance, we study 1D, 2D, 5D, and 10D inputs for two network architectures: Transformer Müller et al. (2022) and CNN. Performance is measured using mean squared error (MSE) and validation loss at convergence (Final Val Loss), summarized in Table 2. GP results are also included as a baseline for MSE. Each backbone is trained with both VA and the proposed DVA. The results show that attention choice has a larger effect than backbone choice. For instance, at 5D, switching a Transformer from VA to DVA reduces MSE from $2.43\!\times\!10^{-4}$ to $2.84\!\times\!10^{-5}$, closer to the GP baseline ($3.42\!\times\!10^{-6}$). The validation loss also improves from 2.04-2.04 to 4.05-4.05—an absolute gain of $2.01$ ($\approx 98.5\%$ relative improvement)—while the CNN-Transformer spread under VA is only $0.25$ ($\approx 10.9\%$). Similarly, at 10D, CNN and Transformer MSEs drop by nearly an order of magnitude under DVA, far exceeding the architecture gap under VA. These results indicate that CNN- and Transformer-based PFNs perform comparably once the attention mechanism is specified, with DVA further pushing performance toward GP quality in higher dimensions.
@@
Comparison with GP: In line with the observations made by authors in Müller et al. (2022), our experiments show that PFNs achieve performance comparable to exact GP inference. As seen in Table 2, PFNs with the proposed DVA consistently move closer to GP performance than those with VA—for instance, in the 10D setting, DVA reduces the MSE from $3.55\text{e-03}$ (CNN-VA) to $5.49\text{e-04}$, compared to the GP baseline of $3.47\text{e-04}$. Importantly, the performance differences between architectures (CNN and Transformer) are relatively minor compared to the gains achieved by changing the attention mechanism, further reinforcing the hypothesis that the effect of attention on PFN performance is far greater than architecture.
We also evaluated the behavior of PFN inference as a function of $N_{\text{context}}$, i.e., How PFN performance improves as the number of available samples increases at inference time?. As shown in Figure 8, PFNs (both CNN and Transformer-based) with the proposed DVA exhibit a consistent decrease in error with increasing context, closely matching the performance of exact GPs in low-dimensional settings. In higher dimensions, GPs maintain a slight advantage, consistent with the trends observed in the training performance analysis. It is important to note that the observed performance gap between 5D and 10D (for both PFNs and GP) arises largely because of limited samples per dataset for 10D model (400 for 5D and 500 for 10D). Similar plots for 1D, 2D PFNs MSE, along with MAE and maximum error for all dimension PFNs are given in the Appendix D.
Figure 3: Comparison with GP: MSE for 5D and 10D PFNs as a function of context size. All models are tested using $n_{\text{test}}=500$, for $N_{context}$. The results show that error consistently decreases with larger context sizes, and that CNN- and Transformer-based PFNs with DVA approach the performance of exact GP inference even in higher dimensions. Exact GP baselines were fit using scikit-learn with $N_{context}$ training samples.

### v2

@@ 2.3 Limitations of Existing PFN Architectures with Joint Attentions
PFNs offer a promising framework for amortized Bayesian inference, though their application to high-dimensional regression has so far been relatively limited hollmann2025tabpfn; wang25d. The common recipe of using a Transformer backbone that performs self-attention over joint $(\mathbf{x},\mathbf{y})$ embeddings has a scaling issue. The design choice of representing each training example $(x,y)$ in PFNs as a joint embedding $\mathrm{enc}(x)+\mathrm{enc}(y)$ can be traced back to the way attention-based models have historically treated their basic units of computation. In natural language processing, the Transformer architecture vaswani2017attention encodes each token as a self-contained representation, i.e. “token as full carrier of information”. In this lineage, PFNs adopt the same strategy by concatenating or summing input and output encodings to form a single token while removing positional encoding müller2024transformersbayesianinference. Below, we examine the this PFN recipe’s structural limitations.
@@
Further, we can also analyze this dimensionality limitation from a bias perspective as the joint embedding breaks localization. The Transformer computes similarity (via dot-product attention) between queries and these mixed embeddings in which the label $\mathbf{y}_{i}$ contributes equally. This conflicts with theoretical results nagler23a, which show that only local samples should influence posterior estimates. Consequently, incorporating joint input–output attention introduces additional bias, which becomes more pronounced in higher dimensions due to the concentration of pairwise distances. In view of these limitations, we propose a simple decoupled value attention (DVA) which keeps the localization intact.
@@ 3 Decoupled-Value Attention
We propose Decoupled-Value Attention (DVA), an input-localized attention mechanism for training PFNs. The proposed DVA is structurally aligned with GP inference by treating input $\mathbf{x}$ and output $\mathbf{y}$ separately at the attention stage. We enforce a strict separation of roles: attention affinities (queries and keys) are computed solely from the inputs, while the aggregated information (values) comes from the corresponding outputs– during both PFN training and prediction. Below, we explain DVA mathematically along with comparative assessment against Vanilla Attention (VA) müller2024transformersbayesianinference and a kernel-based attention wang2025crosskernel.
@@
Training: During training we simulate inference by masking one or more labels from the dataset müller2024transformersbayesianinference. The unmasked pairs $\mathcal{D}_{\mathrm{cx}}=\{(x_{i},y_{i})\}_{i=1}^{N_{\mathrm{context}}}$ form the context set, while the masked inputs $X_{\mathrm{te}}=\{x_{j}\}_{j=1}^{M}$ form the queries. From the context we build $K_{\mathrm{tr}}=W_{k}\varphi_{x}(X_{\mathrm{cx}}),\quad V_{\mathrm{tr}}=W_{v}\varphi_{y}(\mathbf{y}_{\mathrm{cx}})$. Here, $X_{\mathrm{cx}}$ and $\mathbf{y}_{\mathrm{cx}}$ are matrix and vector forms of $\mathcal{D}_{\mathrm{cx}}$ respectively. Further, from the test (masked) inputs query $Q_{\mathrm{te}}$ in matrix from and labels are predicted by attending $H_{\mathrm{te}}$ to the context as :
@@
A head $g(\cdot)$ maps $H_{\mathrm{te}}$ to a predictive distribution, and training minimizes the NLL (equation 1) of the true labels to learn parameters of the network as explained in müller2024transformersbayesianinference.
@@
Vanilla attention is taken from PFN literature müller2024transformersbayesianinference; hollmann2022pfn
@@ 3.1 Localization Effect of DVA and Alignment with GP inference
From equation 5, it is clear that that affinities are determined entirely via relationship between the test input $\mathbf{x}_{\star}$ and context inputs $X_{i}$. Unlike joint embeddings $\phi(x,y)$ in VA, the labels $y_{i}$ do not enter into the similarity measure and only appear downstream through the values as in equation 2. This separation implies that the weight placed on a output $y_{i}$ depends solely on how well $X_{i}$ aligns with $\mathbf{x}_{\star}$ in the projected input space. Thus, the softmax distribution $\alpha_{i}(\mathbf{x}_{\star})$ concentrates mass on a neighborhood of $\mathbf{x}_{\star}$ because the projection matrices and encoders are trained to align nearby inputs with high value of inner product and push apart distant inputs. Consequently, altering labels (outputs) attached to distant inputs cannot affect the prediction asymptotically, which is exactly the localization property required in Theorem 5.4 of nagler23a. This localization property implies that attention weights $\alpha_{i}(\mathbf{x}_{\star})$ associated with context points $\mathbf{x}_{i}$ that are far from the query $\mathbf{x}_{\star}$ should contribute negligibly. Below, we first present a theoretical result which proves that linear embedding-based DVA provides attention weights proportional to Mahalanobis RBF kernel using distance between query and context inputs. Then we also show that under linear separability assumptions, nonlinear embedding DVA also exhibit localization behavior. Thus, connecting the proposed DVA mechanism to the formal notion of localization discussed in nagler23a.
Theorem 1 (DVA attention weight $\propto$ Mahalanobis RBF kernel under linear embeddings)
Assume the input encoder is linear, i.e. $\varphi_{x}(\mathbf{x})=W_{x}\mathbf{x}$ and the DVA query/key maps are $Q=W_{q}W_{x}\mathbf{x},\;K=W_{k}W_{x}\mathbf{x}.$ Let $A:=(W_{q}W_{x})^{\top}(W_{k}W_{x})$. If $A$ is symmetric positive definite, and define $\|\mathbf{z}\|_{A}=\mathbf{z}^{T}A\mathbf{z}$, then for any query $\mathbf{x}_{\star}$ and context point $\mathbf{x}_{i}$ the DVA attention weight equation 5 is proportional to a Mahalanobis RBF kernel
$\alpha_{i}(\mathbf{x}_{\star})\;\propto\;\exp\!\left(-\frac{1}{2\tau}\,\|x_{\star}-x_{i}\|_{A}^{2}\right).$
Theorem 2 (DVA localization under nonlinear embeddings)
Consider a DVA-PFN with fixed query $\mathbf{x}_{\star}$ and context inputs $\mathbf{x}_{i}\in\mathcal{D}_{\texttt{context}}$ of size $N_{\texttt{context}}$. Suppose the input encoder $\varphi_{x}(\cdot)$ is Lipschitz continuous, and that inner product of query $Q_{\star}$ and key $K_{i}$ is locally separable around $\mathbf{x}_{\star}$. Then, as $N_{\texttt{context}}\to\infty$, the total attention mass assigned to all far-away context inputs vanishes with high probability, i.e.: $\sum_{i\in F_{\varepsilon}}\alpha_{i}(\mathbf{x}_{\star})\;\xrightarrow[N_{\texttt{context}}\to\infty]{\mathbb{P}}\;0.$ Where, $F_{\varepsilon}=\{\,i:\|\mathbf{x}_{i}-\mathbf{x}_{\star}\|>\varepsilon\,\}$ denotes the set of “far” indices for any fixed $\varepsilon>0$.
Detailed proofs of both these theorems are provided in Appendix B. Beyond the theoretical guarantees, we also provide empirical evidence that, for DVA-based PFNs, the attention weights $\alpha_{i}$ decay exponentially as the Euclidean distance between the query and context inputs increases. In contrast, VA-based PFNs do not exhibit this behavior (see Fig. 4 and Fig. 5).
We now discuss how DVA aligns with GP inference. As discussed in Section 2.1, GPs model all possible function realizations as zero-mean Gaussians with covariance defined by a kernel, i.e., $f\sim\mathcal{N}(\mathbf{0},K(X,X^{\prime}))$, where $X$ is the input and $K(\cdot,\cdot)$ is the kernel matrix over input pairs. Note that parameterization of the possible function family only depends on the input. Further, for a given kernel hyperparameters, the mean prediction $\mu(\cdot)$ of GP is given as a weighted sum of training dataset outputs with weights solely depending upon inputs williams2006gaussian:
@@
Following equation 4, 5 and 6, the attention weights in DVA can be interpreted as normalized kernel weights that depend only on the inputs. As in kernel smoothing tsai2019transformer, the exponential inner product in $\alpha(\cdot)$ of equation 5 acts as a positive kernel on the input space, with effective bandwidth governed by the scale of the projections and the $1/\sqrt{d_{k}}$ factor. Thus, similar to GP mean prediction, DVA predictions are obtained as weighted sums of training outputs where the weights are determined entirely by input similarity. Readers can refer to tsai2019transformer for more discussion on the relationship between kernel and attention mechanism.
Here, we want to highlight that DVA’s softmax produces non-negative, normalized weights ($\sum\alpha_{i}=1$), whereas the GP coefficients $\beta_{i}(\cdot)$ have no positivity constraint. This limitation is mitigated by subsequent PFN layers (e.g., the final head $g(\cdot)$) and by encoding outputs in the value $V$, which together help adjust the DVA output toward the true GP posterior mean(see Appendix F for effect of final head and value encoder). This construction shows that DVA’s architecture implements a predictor of the form “input-only similarities produce weights, which combine label-dependent values,” precisely matching the dependency structure of a GP.
Another attention choice for PFNs can be kernel-inspired attentions, which relate GP mean weights $\beta(\cdot)$ and PFN attention weights $\alpha(\cdot)$ more closely– while maintaining input localization by decoupling input and output as in DVA. However, if the input affinities are forced through a fixed kernel function, the PFN will become kernel dependent. As discussed before, identifying the best performing kernel is non-trivial and often requires tailoring kernels to specific function classes 9734745. Therefore, it is not advisable to hard-wire a particular kernel in PFN design.
To test the effect of kernel dependence on PFN performance, we design a simple Gaussian kernel (radial basis function, RBF) similarity for attention williams2006gaussian. We emphasize that this formulation is not equivalent to exact GP kernel regression but rather introduces RBF-style distance-based affinities in place of dot-product similarities choromanski2022rethinkingattentionperformers; shen2021efficient. The kernel-based attention assigned to a query input $\mathbf{x}_{\star}$ is then given by
@@
This distance-based attention in equation 7 is more aligned to the RBF kernel; however, it loses flexibility to learn input-localization via training. Thus, the model inherits kernel and $\gamma$ dependence, which may not be suitable for a broader class of functions. To validate this limitation of the kernel-based attention, we test PFN performance with both DVA and this attention. We attempt to learn two classes of functions with different levels of smoothness as shown in Figure 1. The results demonstrate that while kernel-based attention can match DVA in effectively learning smooth functions aligned with the RBF kernel, it under performs on non-smooth functions generated using the linear-periodic kernel. Additionally, we favor DVA over Kernel attention for its superior computational efficiency. By exploiting structural sparsity and hardware-optimized dot products, DVA reduces the computational cost upto four times of dense Kernel attention. We also observed that for RBF training priors, Kernel attention matches performance of DVA while PFN training per epoch is at least four times slower (See Appendix E) for more.
@@ 4 Numerical Results and Discussion
In this section, we present numerical experiments demonstrating the behavior of PFNs equipped with the proposed DVA and with CNN backbone. The results show that with DVA, PFNs a) train with lower validation loss or residual bias, b) CNN, RNN, LSTM and Transformer perform comparably as architecture, underscoring that attention governs training behavior more than backbone architecture and c) remain scalable for learning in complex physical systems. These findings provide empirical support for the theoretical arguments in Section 3.1. Complete experimental details, including architecture choices, hyperparameter selection, and data generation procedures, are provided in the Appendix C, while additional results are provided in Appendix D.
4.1 Bias Reduction and Backbone Architecture Agnostic PFNs
@@
Figure 2: Bias Reduction in PFN Training: Validation loss (NLL) behavior with number of training points for various PFNs (Number of training points = epochs $\times$ steps per epoch $\times$ batch-size $\times$ dataset size. Dataset size is 100 for 1D/2D, 400 for 5D and 500 for 10D PFN). Validation loss was calculated on 64 out-of-sample datasets and Transformer + VA is taken from müller2024transformersbayesianinference.
Backbone Architecture Agnostic PFNs: To analyze the effect of backbone architecture on PFN performance, we study 1D, 2D, 5D, and 10D inputs for two network architectures: Transformer müller2024transformersbayesianinference and CNN along with RNN and LSTM backbones on 1D and 10D input spaces. Performance is measured using mean squared error (MSE) and validation loss at convergence (Final Val Loss), summarized in Table 2. GP results are also included as a baseline for MSE. Each backbone is trained with both VA and the proposed DVA. The results show that attention choice has a larger effect than backbone choice. For instance, at 5D, switching a Transformer from VA to DVA reduces MSE from $2.43\!\times\!10^{-4}$ to $2.84\!\times\!10^{-5}$, closer to the GP baseline ($3.42\!\times\!10^{-6}$). The validation loss also improves from 2.04-2.04 to 4.05-4.05—an absolute gain of $2.01$ ($\approx 98.5\%$ relative improvement)—while the CNN-Transformer spread under VA is only $0.25$ ($\approx 10.9\%$). Similarly, at 10D, CNN and Transformer MSEs drop by nearly an order of magnitude under DVA, far exceeding the architecture gap under VA. These results indicate that CNN- and Transformer-based PFNs perform comparably once the attention mechanism is specified, with DVA further pushing performance toward GP quality in higher dimensions. Additionally, Figure 3 shows that LSTM and RNN backbones can be trained successfully as PFNs, and DVA consistently outperforms vanilla attention in 10D settings across all architectures, while matching the performance in 1D.
Figure 3: Comparison of validation loss vs. training points ($N_{train}$) for RNN and LSTM architectures with VA and DVA attentions.
@@
Comparison with GP: In line with the observations made by authors in müller2024transformersbayesianinference, our experiments show that PFNs achieve performance comparable to exact GP inference. As seen in Table 2, PFNs with the proposed DVA consistently move closer to GP performance than those with VA—for instance, in the 10D setting, DVA reduces the MSE from $3.55\text{e-03}$ (CNN-VA) to $5.49\text{e-04}$, compared to the GP baseline of $3.47\text{e-04}$. Importantly, the performance differences between architectures (CNN and Transformer) are relatively minor compared to the gains achieved by changing the attention mechanism, further reinforcing the hypothesis that the effect of attention on PFN performance is far greater than architecture.
We also evaluated the behavior of PFN inference as a function of $N_{\text{context}}$, i.e., How PFN performance improves as the number of available samples increases at inference time?. As shown in Figure 8, PFNs (both CNN and Transformer-based) with the proposed DVA exhibit a consistent decrease in error with increasing context, closely matching the performance of exact GPs in low-dimensional settings. In higher dimensions, GPs maintain a slight advantage, consistent with the trends observed in the training performance analysis. It is important to note that the observed performance gap between 5D and 10D (for both PFNs and GP) arises largely because of limited samples per dataset for 10D model (400 for 5D and 500 for 10D). Similar plots for 1D, 2D PFNs MSE, along with MAE and maximum error for all-dimensional PFNs, are given in the Appendix D.
To test DVA’s robustness beyond a single function class, we trained and evaluated both DVA and VA on three distinct GP priors generating Smooth, Wiggly (high-frequency), and Mixed complexity functions from RBF kernel. Table 3 shows that DVA consistently outperforms VA across all priors. This confirms DVA’s architecture is robust and not over-tuned to a specific prior. The higher MSEs for the Wiggly and Mixed priors are expected, as their functional complexity presents an intrinsically more challenging learning task for a fixed context size and number of training points.
Uncertainty Quantification & Post-Hoc Localization: In the Appendix H we preovide results indicating that proposed DVA PFN can provide calibrated predictive uncertainty distributions while in Appendix K we compare post-hoc localization nagler23a with the proposed method. Results indicate that in low-dimensions post-hoc localization reduces error and can even slightly outperform DVA. However, in higher-dimensions where VA PFN doesn’t get trained, localization provides no benefit. Thus, emphasizing the need of DVA’s architectural locality.
Table 3: In-distribution performance of DVA vs. VA across diverse GP priors ($N_{\text{context}}=80$).
Training Prior DVA (MSE) VA (MSE) Prior Details ($\ell$: length-scale)
Smooth 1.90e-04 3.16e-04 Fixed lengthscale: $\ell=0.25$
Wiggly 4.19e-03 4.26e-02 Fixed lengthscale: $\ell=0.03$
Mixed 1.39e-02 1.71e-02 Sum of two kernels with $\ell_{1}\sim U(0.1,0.5)$; $\ell_{2}\sim U(0.01,0.04)$
All 2.99e-03 2.43e-02 Per-batch random sampling from the priors above